I don't want the answer I want the procedure using Ti 84 plus to the following question: Compute the standardized test statistic, CHI SQUARED, to test the claim Η2>15.2 if n = 18, s2=21.6, and alpha 0.01
Answer #1

Solution:-

State the hypotheses. The first step is to state the null hypothesis and an alternative hypothesis.

Null hypothesis H0: σ2< 15.2

Alternative hypothesis HA: σ2 > 15.2

Formulate an analysis plan. For this analysis, the significance level is 0.05.

Analyze sample data. Using sample data, the degrees of freedom (DF), and the test statistic (X2).

DF = n - 1 = 18 -1

D.F = 17



We use the Chi-Square Distribution Calculator to find P(Χ2 > 24.158) = 0.115

Steps for finding p-value in Ti 84plus

Step 1) Click on 2nd > VARS > DISTR

Step 2) Select X2CDF

Step 3) Fill lower = 0, Upper = 24.158, D.F = 17

Step 4) Enter Paste and Enter again

Step 5) We get lower cumulative P-value = 0.8848

Step 6) The required p-value = 1 - 0.8848 = 0.1152

Interpret results. Since the P-value (0.1152) is greater than the significance level (0.05), we have to accept the null hypothesis.

We do not have sufficient evidence in the favor of the claim that Η2 >15.2.